# Contracting with a Mechanism Designer

### v1

@@ Appendix C Missing Proofs
$\left\langle\bm{1}-\bm{t}^{\star},\bm{\gamma}\right\rangle\cdot\left(1-F^{n}_{\bm{t}^{\star}}(\theta^{(n)}_{\bm{t}^{\star}})\right)=\left\langle\bm{1}-\alpha^{\star}\cdot\bm{1},\bm{\gamma}\right\rangle\cdot\left(1-F^{n}_{\alpha^{\star}}(\theta^{(n)}_{\alpha^{\star}})\right).$
We can obtain that:
$U_{\mathrm{posted}}$ ${}=\max_{\alpha\in[0,1]}r\left(1-\alpha\right)\cdot\left(1-F^{n}_{\alpha}\left(\theta^{(n)}_{\alpha}\right)\right)$
=${}\overset{(a)}{=}\max_{\alpha\in[0,1]}r\left(1-\alpha\right)\cdot\left(1-F^{n}\left(r\left(1-\alpha\right)+\theta^{(n)}_{\alpha}\right)\right)$
=${}\overset{(b)}{=}\max_{\alpha\in[0,1]}\left(z-\theta^{(n)}_{\alpha}\right)\cdot\left(1-F^{n}\left(z\right)\right)$
=${}\overset{(c)}{=}\max_{\alpha\in[0,1]}\phi^{(n)}(z_{\alpha})\left(1-F^{n}(z_{\alpha})\right)$
=${}\overset{(d)}{=}\max_{z\in\mathtt{supp}(F)}\phi^{(n)}(z)\left(1-F^{n}(z)\right),$
where equality (a) holds by the fact that $F_{\alpha}\left(\theta^{(n)}_{\alpha}\right)=F\left(r\left(1-\alpha\right)+\theta^{(n)}_{\alpha}\right)$, equality (b) holds by letting $z_{\alpha}=r\left(1-\alpha\right)+\theta^{(n)}_{\alpha}$ and equality (c) holds as follows:
$\phi^{(n)}(z_{\alpha})=z_{\alpha}-\frac{1-F^{n}(z_{\alpha})}{nF^{n-1}(z_{\alpha})F^{\prime}(z_{\alpha})}=r(1-\alpha)+\theta^{(n)}_{\alpha}-\frac{1-F^{n}_{\alpha}\left(\theta^{(n)}_{\alpha}\right)}{nF^{n-1}_{\alpha}(z_{\alpha})F^{\prime}_{\alpha,i}\left(\theta^{(n)}_{\alpha}\right)}=r(1-\alpha),$
where the first equality holds by definition, the second equality holds because $z_{\alpha}=r\left(1-\alpha\right)+\theta^{(n)}_{\alpha}$, $F_{\alpha}\left(\theta^{(n)}_{\alpha}\right)=F(z_{\alpha})$ and $F^{\prime}_{\alpha}\left(\theta^{(n)}_{\alpha}\right)=F^{\prime}(z_{\alpha})$, and the last equality is by the the fact that the reserve price $\theta^{(n)}_{\alpha}$ is the zero point of virtual value function $\phi^{(n)}_{\alpha}(\cdot)$.
Finally, we prove equality (d). By Lemma 4.2, the function $\theta_{\alpha}$ (w.r.t. $\alpha$) is continuous in $[0,1]$. Thus, the function $\theta^{(n)}_{\alpha}$ (w.r.t. $\alpha$) is continuous in $[0,1]$. This implies the function $z_{\alpha}$ (w.r.t. $\alpha$) is also continuous in $[0,1]$. Notice that $z_{0}=\theta^{(n)}$ and $z_{1}=r$, where $\theta^{(n)}$ is the monopoly reserve price of distribution $F^{n}$. We have $[\theta^{(n)},r]\subseteq\{z_{\alpha}:\alpha\in[0,1]\}\subseteq\mathtt{supp}(F)$. Hence, equality (c) holds with $\leq$”. On the flip hand, since the contribution distribution $F$ is regular, by the definition of reserve price $\theta^{(n)}$, we have $\phi^{(n)}(z_{\alpha})\leq 0$ if $z_{\alpha}\leq\theta^{(n)}$. Thus, $\phi^{(n)}(z_{\alpha})(1-F^{n}(z_{\alpha}))$ only takes non-negative values if $z_{\alpha}\in[\theta,r]$, leading that equality (c) holds with $\geq$”. Consequently, we have proved equality (c). ∎
C.5 Proof of Theorem 6.2
See 6.2
@@
Lower bounding PoDM and PoA.  Now we analyze the lower bound part of the two ratios, considering Example 5.5 as well. Note that for a single agent (i.e, $n=1$), we have $U^{*}=U_{\mathrm{posted}}$. Therefore, Example 5.5 still serves as the example of a lower bound for the two ratios and the analysis is identical to that of Theorem 5.1. ∎
C.6 Proof of Theorem 6.3
See 6.3
@@
We first show the upper bound parts and then show the lower bound part by analyzing Example 5.7 with a single agent.
Upper bounding PoDM.  We first show the upper bound of PoDM. By Lemma 5.10 we have
$U^{*}\geq\max_{\alpha\in[0,1]}r(1-\alpha)(1-F_{\alpha}^{n}(\theta_{\alpha}^{(n)})))=U_{\mathrm{posted}}\geq\frac{1}{{\mathrm{e}}}\theta^{(n)}(1-F^{n}(\theta^{(n)})).$
@@
$\mathrm{PoDM}$ ${}=\frac{U_{\textsc{sb}}}{U_{\mathrm{posted}}}$
${}\leq{\mathrm{e}}\cdot\frac{U_{\textsc{sb}}}{\theta^{(n)}(1-F^{n}(\theta^{(n)}))}$
=${}\overset{(a)}{=}{\mathrm{e}}\cdot 1.2683$
${}\leq 3.448,$
where equality (a) holds by Lemma 5.12.
@@
Lower bounding PoDM and PoA.  Now we analyze the lower bound part of the two ratios by considering Example 5.7 with only a single agent (i.e, $n=1$). Since for a single agent, we have $U^{*}=U_{\mathrm{posted}}$, Example 5.7 still serves as the example of a lower bound for the two ratios and the analysis is identical to that of Theorem 5.6, where we assign the number of agent $n=1$. ∎
C.7 Proof of Theorem 7.3
@@
Next, we will prove Theorem 7.3, which provide negative results for the case that the principal has little information about the market size.
See 7.3
@@
Notice that the monopoly reserve price $\theta_{\alpha}$ is positive. By definition, given any $\alpha\in[0,1]$, we have that
@@
[…]
@@
[…]

### v2

@@ Appendix C Missing Proofs
$\left\langle\bm{1}-\bm{t}^{\star},\bm{\gamma}\right\rangle\cdot\left(1-F^{n}_{\bm{t}^{\star}}(\theta^{(n)}_{\bm{t}^{\star}})\right)=\left\langle\bm{1}-\alpha^{\star}\cdot\bm{1},\bm{\gamma}\right\rangle\cdot\left(1-F^{n}_{\alpha^{\star}}(\theta^{(n)}_{\alpha^{\star}})\right)$
which completes the proof as desired. ∎
C.7 Proof of Theorem 5.2
In this section, we prove Theorem 5.2.
See 5.2
@@
Lower bounding PoDM and PoA. Now we analyze the lower bound part of the two ratios, considering Example C.2 as well. Note that for a single agent (i.e, $n=1$), we have $U^{*}=U_{\mathrm{posted}}$. Therefore, Example C.2 still serves as the example of a lower bound for the two ratios and the analysis is identical to that of Theorem 4.1. ∎
C.8 Proof of Theorem 5.3
In this section, we prove Theorem 5.3.
See 5.3
@@
We first show the upper bound parts and then show the lower bound part by analyzing Example C.4 with a single agent.
Upper bounding PoDM. We first show the upper bound of PoDM. By Lemma C.12 we have
$U^{*}\geq\max_{\alpha\in[0,1]}r(1-\alpha)(1-F_{\alpha}^{n}(\theta_{\alpha}^{(n)})))=U_{\mathrm{posted}}\geq\frac{1}{{\mathrm{e}}}\cdot\theta^{(n)}(1-F^{n}(\theta^{(n)})).$
@@
$\mathrm{PoDM}{}=\frac{U_{\textsc{sb}}}{U_{\mathrm{posted}}}{}\leq{\mathrm{e}}\cdot\frac{U_{\textsc{sb}}}{\theta^{(n)}(1-F^{n}(\theta^{(n)}))}{}\overset{(a)}{=}{\mathrm{e}}\cdot 1.2683{}\leq 3.448,$
where equality (a) holds by Lemma C.14.
@@
Lower bounding PoDM and PoA. Now we analyze the lower bound part of the two ratios by considering Example C.4 with only a single agent (i.e., $n=1$). Since for a single agent, we have $U^{*}=U_{\mathrm{posted}}$, Example C.4 still serves as the example of a lower bound for the two ratios. Therefore, the analysis is identical to that of Theorem 4.2, where we assign $n=1$. ∎
C.9 Proof of Theorem 6.1
In this section, we prove Theorem 6.1.
See 6.1
Our analysis relies on the following technical lemmas.
Lemma C.17.
Let $s,\ell\in\mathbb{N}$ with $s\leq\ell$. Then, for any regular distribution $F$ and contract $\alpha\in[0,1]$, we have
$\frac{U(\alpha,s)}{U(\alpha,\ell)}\geq\frac{s}{\ell}.$
Proof.
Recall that $U(\alpha,n)=\phi(z)(1-F^{n}(z))$, where $z=r(1-\alpha)+\theta_{\alpha}$. We have
$\frac{U(\alpha,s)}{U(\alpha,\ell)}=\frac{\phi(z)(1-F^{s}(z))}{\phi(z)(1-F^{\ell}(z))}={}\frac{1-F^{s}(z)}{1-F^{\ell}(z)}.$
Consider the function $x\mapsto\frac{1-x^{s}}{1-x^{\ell}}$ for $x\in[0,1]$. One can easily verify that this function is decreasing for $x\in[0,1]$. Thus, we have
$\frac{1-F^{s}(z)}{1-F^{\ell}(z)}\geq\lim_{x\to 1}\frac{1-x^{s}}{1-x^{\ell}}=\lim_{x\to 1}\frac{-sx^{s-1}}{-\ell x^{\ell-1}}=\frac{s}{\ell},$
which gives the desired inequality. ∎
Lemma C.18.
Let $s,\ell\in\mathbb{N}$ with $s\leq\ell$. Then, for any regular distribution $F$, we have
$\frac{U_{\textsc{fb}}(s)}{U_{\textsc{fb}}(\ell)}\geq\frac{U_{\textsc{sb}}(s)}{U_{\textsc{sb}}(\ell)}\geq\frac{s}{\ell}.$
Proof.
We first prove the second inequality. For any $n\in\mathbb{R}^{+}$, we can rewrite $U_{\textsc{fb}}(n)$ as
$U_{\textsc{sb}}(n)=\int_{0}^{\infty}\phi(z)(1-F^{n}(z))\cdot{\mathrm{d}}z=\int_{0}^{\infty}\phi(z)\int_{0}^{n}F^{u}(z)(-\ln F(z))\cdot{\mathrm{d}}u{\mathrm{d}}z.$
Since the above integration is non-negative and measurable, according to Tonelli’s theorem, we can exchange the order of integration as
$U_{\textsc{sb}}(n)=\int_{0}^{n}\int_{0}^{\infty}\phi(z)F^{u}(z)(-\ln F(z))\cdot{\mathrm{d}}z{\mathrm{d}}u.$
Define the function $g(u)=\int_{0}^{\infty}\phi(z)F^{u}(z)(-\ln F(z))\cdot{\mathrm{d}}z$. We know $U_{\textsc{sb}}(s)=\int_{0}^{s}g(u)\cdot{\mathrm{d}}u$ and $U_{\textsc{sb}}(\ell)=\int_{0}^{\ell}g(u)\cdot{\mathrm{d}}u$. Thus, the desired inequality is equivalent to
$\frac{1}{s}\int_{0}^{s}g(u)\cdot{\mathrm{d}}u\geq\frac{1}{\ell}\int_{0}^{\ell}g(u)\cdot{\mathrm{d}}u.$
Therefore, it suffices to show that $g(\cdot)$ is non-increasing according to the monotonicity of the average integral.
Fix $u_{1},u_{2}\in\mathbb{R}^{+}$ with $u_{1}\leq u_{2}$. To show that $g(u_{2})\leq g(u_{1})$, we employ the integral Chebyshev inequality: for a measure $\mu$ and two functions $\xi$ and $\zeta$ with opposite monotonicity. it holds
$\int{\mathrm{d}}\mu\int\xi\zeta\cdot{\mathrm{d}}\mu\leq\int\xi\cdot{\mathrm{d}}\mu\int\zeta\cdot{\mathrm{d}}\mu.$
Let ${\mathrm{d}}\mu(z)=F^{u_{1}}(z)\cdot{\mathrm{d}}z$. Define $\xi(z)=\phi(z)(-\ln F(z))$ and $\zeta(z)=F^{u_{2}-u_{1}}(z)$. Given that $F$ is regular, $\xi(\cdot)$ is non-increasing, while $\zeta(\cdot)$ is non-decreasing. Applying the inequality yields
$g(u_{2})={}$ $\int_{0}^{\infty}\xi(z)\zeta(z)\cdot{\mathrm{d}}\mu(z)$
$\leq{}$ $\frac{1}{\int_{0}^{\infty}{\mathrm{d}}\mu(z)}\int_{0}^{\infty}\xi(z)\cdot{\mathrm{d}}\mu(z)\int_{0}^{\infty}\zeta(z)\cdot{\mathrm{d}}\mu(z)$
$={}$ $g(u_{1})\cdot\frac{\int_{0}^{\infty}F^{u_{2}}(z)\cdot{\mathrm{d}}z}{\int_{0}^{\infty}F^{u_{1}}(z)\cdot{\mathrm{d}}z}$
$\leq{}$ $g(u_{1}),$
where the last inequality holds since $F^{u}(z)$ is non-increasing in $u$ for any fixed $z$. Consequently, $g(u)$ is non-increasing for $u\in\mathbb{R}^{+}$, implying the desired inequality.
Next, we prove the first inequality. By definition, it is sufficient to prove that
$\frac{\int_{0}^{\infty}(1-F^{s}(z))\cdot{\mathrm{d}}z}{\int_{0}^{\infty}(1-F^{\ell}(z))\cdot{\mathrm{d}}z}\geq\frac{\int_{0}^{\infty}\phi(z)(1-F^{s}(z))\cdot{\mathrm{d}}z}{\int_{0}^{\infty}\phi(z)(1-F^{\ell}(z))\cdot{\mathrm{d}}z}.$
We can rearrange the above inequality as
$\int_{0}^{\infty}\phi(z)(1-F^{\ell}(z))\cdot{\mathrm{d}}z\int_{0}^{\infty}(1-F^{s}(z^{{\dagger}}))\cdot{\mathrm{d}}z^{{\dagger}}\geq\int_{0}^{\infty}\phi(z^{{\dagger}})(1-F^{s}(z^{{\dagger}}))\cdot{\mathrm{d}}z^{{\dagger}}\int_{0}^{\infty}(1-F^{\ell}(z))\cdot{\mathrm{d}}z,$
which is equivalent to
$\int_{0}^{\infty}\int_{0}^{\infty}\left(\phi(z)-\phi(z^{{\dagger}})\right)(1-F^{\ell}(z))(1-F^{s}(z^{{\dagger}}))\cdot{\mathrm{d}}z{\mathrm{d}}z^{{\dagger}}\geq 0.$
Exchanging the variables $z$ and $z^{{\dagger}}$ in the double integral and taking the average, we can further derive that
$\int_{0}^{\infty}\int_{0}^{\infty}\left(\phi(z)-\phi(z^{{\dagger}})\right)(1-F^{\ell}(z))(1-F^{s}(z^{{\dagger}}))\cdot{\mathrm{d}}z{\mathrm{d}}z^{{\dagger}}$
$={}$ $\frac{1}{2}\int_{0}^{\infty}\int_{0}^{\infty}\left(\phi(z)-\phi(z^{{\dagger}})\right)\left((1-F^{\ell}(z))(1-F^{s}(z^{{\dagger}}))-(1-F^{\ell}(z^{{\dagger}}))(1-F^{s}(z))\right)\cdot{\mathrm{d}}z{\mathrm{d}}z^{{\dagger}}$
$={}$ $\int_{0}^{\infty}\int_{z^{{\dagger}}}^{\infty}\left(\phi(z)-\phi(z^{{\dagger}})\right)\left((1-F^{\ell}(z))(1-F^{s}(z^{{\dagger}}))-(1-F^{\ell}(z^{{\dagger}}))(1-F^{s}(z))\right)\cdot{\mathrm{d}}z{\mathrm{d}}z^{{\dagger}}.$
Since the integrand is symmetric in $z$ and $z^{{\dagger}}$, the second equality holds by splitting the integral into two parts where $z\geq z^{{\dagger}}$ and $z<z^{{\dagger}}$.
Therefore, it is sufficient to show
$\int_{0}^{\infty}\int_{z^{{\dagger}}}^{\infty}\left(\phi(z)-\phi(z^{{\dagger}})\right)A(z,z^{{\dagger}})\geq 0,$
where $A(z,z^{{\dagger}})=(1-F^{\ell}(z))(1-F^{s}(z^{{\dagger}}))-(1-F^{\ell}(z^{{\dagger}}))(1-F^{s}(z))$. One can easily verify that the function $x\mapsto\frac{1-x^{s}}{1-x^{\ell}}$ is decreasing for $x\in[0,1]$. Thus, we know $A(z,z^{{\dagger}})\geq 0$ for any $z\geq z^{{\dagger}}$. In addition, since $F$ is a regular distribution, its virtual value function $\phi(\cdot)$ is non-decreasing. Thus, we have $\phi(z)-\phi(z^{{\dagger}})\geq 0$ for any $z\geq z^{{\dagger}}$. Consequently, the integrand is always non-negative for any $z\geq z^{{\dagger}}$, which proves the desired inequality. ∎
Now, we are ready to prove Theorem 6.1.
Proof of Theorem 6.1..
Recall that the regular contribution distribution $F$ has the bounded support $\mathtt{supp}(F)=[\underline{w},\bar{w}]$, $(0\leq\underline{w}\leq\bar{w}\leq r)$, and the ratio of its upper and lower bounds satisfies $1\leq\bar{w}/\underline{w}\leq\kappa$.
Consider the design of setting $\alpha^{{\dagger}}$ as the optimal contract for some number of agents $n^{{\dagger}}$, i.e.,
$\alpha^{{\dagger}}=\operatorname*{arg\,max}_{\alpha\in[0,1]}U(\alpha,n^{{\dagger}})=\operatorname*{arg\,max}_{z\in\mathtt{supp}(F)}\phi(z)(1-F^{n^{{\dagger}}}(z)).$
Let $n^{*}\in[s:\ell]$ denote the integer that maximizes $\frac{U_{\textsc{fb}}(n)}{U(\alpha^{{\dagger}},n)}$ under this contract $\alpha^{{\dagger}}$, i.e.,
$n^{*}=\operatorname*{arg\,max}_{n\in[s:\ell]}\frac{U_{\textsc{fb}}(n)}{U(\alpha^{{\dagger}},n)}.$
By definition, we have
$\mathrm{PoA}(s,\ell)=\min_{\alpha\in[0,1]}\max_{n\in[s:\ell]}\frac{U_{\textsc{fb}}(n)}{U(\alpha,n)}\leq\max_{n\in[s:\ell]}\frac{U_{\textsc{fb}}(n)}{U(\alpha^{{\dagger}},n)}=\frac{U_{\textsc{fb}}(n^{*})}{U(\alpha^{{\dagger}},n^{*})}.$
Using the monotonicity of $U(\alpha^{{\dagger}},\cdot)$ and $U_{\textsc{fb}}(\cdot)$, we bound this ratio according to the relation between $n^{{\dagger}}$ and $n^{*}$. When $n^{{\dagger}}\leq n^{*}$, by Lemma C.18, we have
$\mathrm{PoA}(s,\ell)\leq{}\frac{U_{\textsc{fb}}(n^{*})}{U(\alpha^{{\dagger}},n^{*})}={}\frac{U_{\textsc{fb}}(n^{*})}{U_{\textsc{fb}}(n^{{\dagger}})}\cdot\frac{U_{\textsc{fb}}(n^{{\dagger}})}{U(\alpha^{{\dagger}},n^{*})}\leq\frac{n^{*}}{n^{{\dagger}}}\cdot\frac{U_{\textsc{fb}}(n^{{\dagger}})}{U(\alpha^{{\dagger}},n^{{\dagger}})}\leq\frac{\ell}{s}\cdot\mathcal{O}_{\kappa}\left((\log\kappa)^{2}\right).$
When $n^{{\dagger}}>n^{*}$, by Lemma C.17, we have
$\mathrm{PoA}(s,\ell)\leq{}\frac{U_{\textsc{fb}}(n^{*})}{U(\alpha^{{\dagger}},n^{*})}={}\frac{U_{\textsc{fb}}(n^{*})}{U(\alpha^{{\dagger}},n^{{\dagger}})}\cdot\frac{U(\alpha^{{\dagger}},n^{{\dagger}})}{U(\alpha^{{\dagger}},n^{*})}\leq\frac{U_{\textsc{fb}}(n^{{\dagger}})}{U(\alpha^{{\dagger}},n^{{\dagger}})}\cdot\frac{n^{{\dagger}}}{n^{*}}\leq\mathcal{O}_{\kappa}\left((\log\kappa)^{2}\right)\cdot\frac{\ell}{s}.$
Similarly, we can also obtain that by Lemma C.17 that
$\mathrm{PoDM}(s,\ell)\leq\max_{n\in[s:\ell]}\frac{U_{\textsc{sb}}(n)}{U(\alpha^{{\dagger}},n)}\leq\frac{\ell}{s}\cdot\mathcal{O}_{\kappa}\left(\log\kappa\right).$
which completes the proof. ∎
C.10 Proof of Theorem 6.2
In this section, we prove Theorem 6.2.
See 6.2
Lemma C.19 (Babaioff et al., 2017, Lemma 5.3).
Let $F$ be a positive MHR distribution. Let $\mu_{n}$ be the expected value of the first-order static out of $n$ i.i.d. variables drawing from distribution $F$. For all $s\leq\ell$, it holds
Our analysis relies on the following technical lemmas.
Lemma C.20.
Proof.
Recall that $\bar{\mu}_{n}$ is the excepted value of the first-order static out of $n$ i.i.d. variables drawing from distribution $\bar{F}$. We also let $\bar{\mu}_{0}=0$. Define t=1F(0)t=1-F(0) and $B(n,k)=\binom{n}{k}t^{k}(1-t)^{n-k}$. Using the binomial theorem as follows, the first-best benchmark $U_{\textsc{fb}}(n)$ can be represented by $\bar{\mu}_{i}$ ($i\in[n]$) as follows.
$={}$ $\int_{0}^{\infty}z\cdot{\mathrm{d}}F^{n}(z)$
$={}$ $\int_{0}^{\infty}z\cdot{\mathrm{d}}(t\bar{F}(z)+(1-t))^{n}$
$={}$ $\int_{0}^{\infty}z\cdot{\mathrm{d}}\sum_{i=0}^{n}B(n,i)\bar{F}^{i}(z)$
$={}$ $\sum_{i=0}^{n}B(n,i)\cdot\int_{0}^{\infty}z\cdot{\mathrm{d}}\bar{F}^{i}(z)$
$={}$ $\sum_{i=0}^{n}B(n,i)\cdot\bar{\mu}_{i}.$
$U_{\textsc{fb}}(n)={}$ $\sum_{i=0}^{n}B(s,i)\mathcal{H}_{i}$
$={}$ $\sum_{i=0}^{n}\binom{n}{i}t^{i}(1-t)^{n-i}\cdot\sum_{j=1}^{i}\binom{i}{j}\frac{(-1)^{j}}{j}$
$={}$ $\sum_{i=0}^{n}\sum_{j=1}^{i}\binom{n}{i}\binom{i}{j}t^{i}(1-t)^{n-i}\cdot\int_{0}^{1}-(-z)^{j-1}\cdot{\mathrm{d}}z$
$={}$ $\int_{0}^{1}-\sum_{i=0}^{n}\sum_{j=1}^{i}\binom{n}{i}\binom{i}{j}t^{i}(1-t)^{n-i}(-z)^{j-1}\cdot{\mathrm{d}}z.$
$={}$ $\int_{0}^{1}\frac{1}{z}\sum_{i=0}^{n}\sum_{j=1}^{i}\binom{n}{i}\binom{i}{j}t^{i}(-(1-t)z)^{j}(1-t)^{n-i-j}\cdot{\mathrm{d}}z$
$={}$ $\int_{0}^{1}\frac{1}{z}\left(\sum_{i=0}^{n}\sum_{j=0}^{i}\binom{n}{i}\binom{i}{j}t^{i}(-(1-t)z)^{j}(1-t)^{n-i-j}-\sum_{i=0}^{n}\binom{n}{i}t^{i}(1-t)^{n-i}\right)\cdot{\mathrm{d}}z$
$=$ $\int_{0}^{1}\frac{(t-(1-t)z+1-t)^{n}-1}{z}\cdot{\mathrm{d}}z$
$=$ $\int_{0}^{1}\frac{(1-(1-t)z)^{n}-1}{z}\cdot{\mathrm{d}}z.$
$U_{\textsc{fb}}(n)=$ $\int_{0}^{1}\frac{(1-(1-t)z)^{n}-1}{z}\cdot{\mathrm{d}}z.$
which completes the proof. ∎
Lemma C.21.
Define $\hat{\mathcal{H}}_{0}=0$ and
$\hat{\mathcal{H}}_{n}=\int_{0}^{1}\frac{1-(tz+1-t)^{n}}{1-z}{\mathrm{d}}z$
$\frac{\hat{\mathcal{H}}_{s}}{\hat{\mathcal{H}}_{\ell}}\geq\frac{\ln(\frac{1}{2}ts+1)}{\ln(\frac{1}{2}t\ell+1)}.$
Proof.
By induction, it is sufficient to show that
$\frac{\hat{\mathcal{H}}_{n+1}}{\hat{\mathcal{H}}_{n}}\leq\frac{\ln(\delta t(n+1)+1)}{\ln(\delta tn+1)},$
for any integer $n\in\mathbb{N}$, where $\delta=\frac{1}{2}$. Letting $\Delta_{n}=\hat{\mathcal{H}}_{n+1}-\hat{\mathcal{H}}_{n}$, it is sufficient to show that
which is equivalent to showing
We rewrite the above condition as
Therefore, by induction, it is sufficient to show that
for any integer $n\in\mathbb{N}^{+}$.
Define a real function $\xi\colon z\mapsto z\ln\left(\frac{1}{z}+1\right)$. We know $\xi(\cdot)$ is increasing since
which is equivalent to the fact that
Finally, according to Lemma C.22 below, we finally derive that
which gives us the desired conclusion. ∎
Lemma C.22.
Define $\hat{\mathcal{H}}_{0}=0$ and
$\hat{\mathcal{H}}_{n}=\int_{0}^{1}\frac{1-(tz+1-t)^{n}}{1-z}\cdot{\mathrm{d}}z,$
$\frac{\hat{\mathcal{H}}_{n}-\hat{\mathcal{H}}_{n-1}}{\hat{\mathcal{H}}_{n+1}-\hat{\mathcal{H}}_{n}}\geq\frac{\frac{1}{2}tn+1}{\frac{1}{2}t(n-1)+1}.$
Proof.
$\frac{\Delta_{n-1}}{\Delta_{n}}\geq\frac{\delta tn+1}{\delta t(n-1)+1}.$
First, by the definition of $\hat{\mathcal{H}}_{n}$, we have
$\Delta_{n}={}$ $\hat{\mathcal{H}}_{n+1}-\hat{\mathcal{H}}_{n}$
$={}$ $\int_{0}^{1}\left(\frac{1-(tz+1-t)^{n+1}}{1-z}-\frac{1-(tz+1-t)^{n}}{1-z}\right)\cdot{\mathrm{d}}z$
$={}$ $t\cdot\int_{0}^{1}(tz+1-t)^{n}\cdot{\mathrm{d}}z$
$={}$ $\frac{1-(1-t)^{n+1}}{n+1}.$
To show the desired inequality, we construct the following difference:
$\Delta_{n-1}\cdot(\delta t(n-1)+1)-\Delta_{n}\cdot(\delta tn+1)$
$={}$ $(1-(1-t)^{n})\cdot\left(\delta t+\frac{1-\delta t}{n}\right)-(1-(1-t)^{n+1})\cdot\left(\delta t+\frac{1-\delta t}{n+1}\right)$
$={}$ $\frac{1-\delta t}{n(n+1)}-(1-t)^{n}\cdot\left(\delta t+\frac{1-\delta t}{n}\right)+(1-t)^{n+1}\cdot\left(\delta t+\frac{1-\delta t}{n+1}\right)$
$={}$ $\frac{1-\delta t}{n(n+1)}-(1-t)^{n}\cdot\left(\delta t^{2}+\frac{(nt+1)(1-\delta t)}{n(n+1)}\right)$
$={}$ $\frac{1-\delta t}{n(n+1)}-(1-t)^{n}\cdot\frac{\delta t^{2}n^{2}+tn+1-\delta t}{n(n+1)}.$
This indicates that we only need to show the function
is non-negative when $z\in\mathbb{N}$. Observe $g_{t}(0)=0$, it is sufficient to show $g_{t}(\cdot)$ is weakly increasing for any fixed $t\in[0,1]$.
Notice that the first derivation of the function $g_{t}(\cdot)$ can be calculated as
Clearly, we have $g^{\prime}_{1}(z)=2\delta z+1\geq 0$, indicating that $g_{1}(\cdot)$ is weakly increasing. Now, we assume that $t\in[0,1)$.
Consider the function $\frac{-g^{\prime}_{t}(z)}{(1-t)^{z}}$ and analysis the sign of this function for any $t\in[0,1)$. On the one hand, the value of this function at the point $z=0$ is
On the other hand, since $\ln(1-t)<0$ for $t\in[0,1)$, the first derivation of this function can be calculated as
which implies that the function $g^{\prime}_{t}(z)/(1-t)^{z}$ is weakly decreasing. Combining the above two facts, we derive that for any $z\geq 0$, it holds $g^{\prime}_{t}(z)/(1-t)^{z}\leq 0$, leading to $g^{\prime}_{t}(z)\geq 0$.
Lemma C.23.
Proof.
Let $\bar{F}$ be the conditional distribution conditioning on the event that the random value following $F$ is non-negative. Denote by $\bar{\mu}_{n}$ the expected value of the first-order statistic out of $n$ i.i.d. variables drawing from distribution $\bar{F}$. Define t=1F(0)t=1-F(0) and $B(n,k)=\binom{n}{k}t^{k}(1-t)^{n-k}$. According to Lemma C.20, we know
In addition, by using the Chernoff inequalities of $B(\cdot,\cdot)$ and Lemma C.19, we further obtain that
$\geq{}$ $\frac{\sum_{i=0}^{s}B(s,i)\mathcal{H}_{i}-(\mathcal{H}_{s}-\mathcal{H}_{k})\sum_{i=\lfloor k\rfloor}^{s}B(s,i)}{\sum_{i=0}^{\ell}B(\ell,i)\mathcal{H}_{i}+\mathcal{H}_{k}\sum_{i=0}^{\lceil k\rceil}B(\ell,i)}$
$\geq{}$ $\frac{\sum_{i=0}^{s}B(s,i)\mathcal{H}_{i}-(\mathcal{H}_{s}-\mathcal{H}_{k}){\mathrm{e}}^{-2s(1-t-(1-\lfloor k\rfloor/n))^{2})}}{\sum_{i=0}^{\ell}B(\ell,i)\mathcal{H}_{i}+\mathcal{H}_{k}{\mathrm{e}}^{-2\ell(t-\lceil k\rceil/n)^{2})}}$
$\geq{}$ $\frac{\sum_{i=0}^{s}B(s,i)\mathcal{H}_{i}}{\sum_{i=0}^{\ell}B(\ell,i)\mathcal{H}_{i}}-\mathcal{O}_{s,\ell}\left(\ln\ell\cdot{\mathrm{e}}^{-t\sqrt{s\ell}}\right).$
Recall that $\hat{\mathcal{H}}_{0}=0$ and
$\hat{\mathcal{H}}_{n}=\int_{0}^{1}\frac{1-(tz+1-t)^{n}}{1-z}\cdot{\mathrm{d}}z,$
for any positive number $n\in\mathbb{N}^{+}$ and real $t\in[0,1]$. Combining lemmas C.20 and C.21, it directly follows that
which completes the proof. ∎
Now, we are ready to prove Theorem 6.2.
Proof of Theorem 6.2..
Suppose that for each agent $i\in[n]$, the contribution $w_{i}$ satisfies an MHR distribution $F$. According to Lemma C.23, we have
Exploiting the monotonicity of $U(\alpha^{{\dagger}},\cdot)$ and $U_{\textsc{fb}}(\cdot)$, $\mathrm{PoA}(s,\ell)$ can be upper bounded by
$\mathrm{PoA}(s,\ell)\leq\max_{n\in[s:\ell]}\frac{U_{\textsc{fb}}(n)}{U(\alpha^{{\dagger}},n)}\leq\frac{U_{\textsc{fb}}(\ell)}{U\left(\alpha^{{\dagger}},s\right)}\leq\frac{U_{\textsc{fb}}(\ell)}{U_{\textsc{fb}}(s)}\cdot\frac{U_{\textsc{fb}}(s)}{U\left(\alpha^{{\dagger}},s\right)}\leq\frac{\ln\left(\frac{1}{2}(1-F(0))\ell+1\right)}{\ln\left(\frac{1}{2}(1-F(0))s+1\right)}\cdot{\mathrm{e}}^{2}+o_{s,\ell}\left(1\right),$
where the last inequality holds by the definition of $\alpha^{{\dagger}}$ and Theorem 4.2. Moreover, by Theorem 4.3, we can also obtain that
which completes the proof. ∎
C.11 Proof of Theorem 6.3
@@
In this section, we will prove Theorem 6.3, which provide negative results for the case that the principal has little information about the market size.
See 6.3
@@
Notice that the monopoly reserve price $\theta_{\alpha}$ is positive. By definition, given any $\alpha\in[0,1]$, we have
@@
$\leq{}$ $r(1-\alpha)\left(1-F_{r}^{n}(r(1-\alpha))\right)$
$={}$ $r(1-\alpha)\left(1-\left(\frac{1-{\mathrm{e}}^{-r(1-\alpha)}}{1-{\mathrm{e}}^{-r}}\right)^{n}\right)$
$\leq{}$ $r(1-\alpha)\left(1-\left(1-{\mathrm{e}}^{-r(1-\alpha)}\right)^{n}\right),$
@@
$\geq{}$ $\int_{1}^{r}(z-1)\cdot{\mathrm{d}}F_{r}^{\ell}(z)$
$={}$ $r-1-\int_{0}^{r-1}F_{r}^{\ell}(z+1)\cdot{\mathrm{d}}z$
$={}$ $r-1-\int_{0}^{r-1}\left(1-\frac{1-{\mathrm{e}}^{-z-1}}{1-{\mathrm{e}}^{-r}}\right)^{\ell}\cdot{\mathrm{d}}z$
$\geq{}$ $r-1-(r-1)\left(1-\frac{1-{\mathrm{e}}^{-1}}{1-{\mathrm{e}}^{-r}}\right)^{\ell}$
$\geq{}$ $r-1-(r-1){\mathrm{e}}^{-\ell}$
$\geq{}$ $r-1-\varepsilon.$